Two capacitor of capacitance 6 MF and 12 m f are connected in series with a battery the voltage across the 6mf capacitor is 2v comput total battery voltage

Grade:12

1 Answers

Adarsh
6 years ago
Dear Guddu,
In series combination the charge on each capacitor remains same as same current flows through it .
Now, charge on 6mf capacitor ,q=c.v=12mC.
same charge will also present on 12mf capacitor so total charge on capacitors= 24mf.
Resultant capacitance= 4mf
Now, V=Q/C = 24/4 = 6V.
Thanks and Regards
Adarsh